a)     A woman who has won a prize is offered a lump sum of 100000 to invest now or 55000 to invest at the end of this year .If all investments are assumed to earn 7% per annum which should she choose if she wants to withdraw the money after

i)                  4 years

2 years             


1
Expert's answer
2022-03-03T09:12:45-0500

A) If she invests the 100000 as a lump sum, after 4 years, she'd be able to withdraw

"FV=100000(1+0.07)^4"

"FV_4=131079.60"


If she invests 55000 at the end of the year, after 4 years, she'd be able to withdraw

"FV=\frac{55000[(1+0.07)^4-1]}{0.07}"

"FV_4=244196.87"


She'd choose to invest the 55000 at the end of the year as it pays more after 4 years.


B) if she chooses the 100000 plan to withdraw after 2 years:

"FV=100000(1+0.07)^2"

"FV_2=114490"


If she chooses the 55000 plan to withdraw after 2 years:

"FV=\frac{55000[(1+0.07)^2-1]}{0.07}"

"FV_2=113850"


She'd choose the 100000 plan to withdraw because it pays more after 2 years.